The multiple regression model represents pricing for residential housing in a certain market. Predicted Price ̂ = 19,856.56 + 6,985.25 bedrooms + 87.53 square foot. A house in the local market has 5 bedrooms and 3,200 square feet of living area. Use the multiple regression model to determine the price and the residual if the house sells for $352,200.

  • A. predicted price = $334,879, error = $17,321
  • B. predicted price = $334,879, error = – $17,700
  • C. predicted price = $355,700, error = – $18,321
  • D. predicted price = $334,879, error = – $20,000
Answer #1

The multiple regression model for prices of residential housing is pre.price=19856.56+6985.25 bedrooms+87.53 squrefoot

Given that, no.of bedrooms=5 , living area=3200 sq.foot then by using above multiple regression model predicted price is,

Price=19856.56+(6985.25 ×5)+(87.53 ×3200)

Predicted price=$ 334879

Error= observed value- predicted value

=352200- 334879

=$ 17321

Hence, answer is option A